# Constraining Participation: Affordances of Feedback Features in Interfaces to Large Language Models

### v1

@@ 4.2.1. Perception
Respondents generally perceived the feedback features in the interface as accessible and user-friendly. Over three-quarters reported confidence in their ability to evaluate and submit feedback about the performance and/or impact of generative AI systems such as ChatGPT (“Strongly Agree” (21.8%) or “Somewhat Agree” (56.5%), n=524). Of those who submitted at least one of the three feedback modes, almost all found doing so through the features extremely or somewhat easy (95.1% n=442).
However, some users were more confident than others in submitting feedback. Younger users (18-34) and those who frequently engaged with AI-related topics expressed more confidence than would be expected if there was no relationship between these factors and confidence. Conversely, users with less frequent engagement were less confident than would be expected. We found weak associations between confidence and age ($\chi^{2}$ = 17.068, adj. $p$-value = 0.045, df = 6, Cramér’s V = 0.131) and frequent engagement with AI-related topics ($\chi^{2}$2 = 20.726, adj. $p$-value = 0.045, df = 8, Cramér’s V = 0.144).
@@ 4.2.2. Dexterity
Few respondents reported barriers to accessing ChatGPT or submitting feedback, suggesting that users with varying levels of dexterity can provide input. When asked about barriers to accessing ChatGPT or submitting feedback, most respondents (77.3%, n=519) reported experiencing no barriers. Among those who encountered barriers, the most common were language and internet connectivity issues. We did not find significant associations between most demographic factors or AI experience and knowledge factors and those who experienced barriers. While those working in AI were more likely to report experiencing barriers than would be expected if there was no relationship between the variables, the association was weak ($\chi^{2}$ = 9.821, adj. $p$-value = 0.023, df = 1, Cramér’s V = 0.143).
However, the requirement to prompt the dialogue agent before submitting feedback means only users can submit feedback through the interface’s features. Our sample, therefore, does not include non-user perspectives, and given that most respondents used ChatGPT at least weekly, it also includes limited input from infrequent users. Notably, of those who completed the survey but did not indicate they had submitted feedback through one of the three options, the most common reason was that they were unaware of the feedback option (34.6%, n=78).
While the ChatGPT interface may constrain direct feedback, respondents found alternative ways to express their opinions and participate in discussions about the system. Most respondents (80.4%, n=526) did not indicate they shared feedback about ChatGPT through other channels or platforms. However, those who did primarily used X/Twitter (42%), followed by Facebook (18%) and Reddit (17%). Several respondents engaged with dedicated communities, including subreddits like r/ChatGPT and online forums like the OpenAI Developer Forum. Most respondents who did indicate they had shared feedback through other channels or platforms (71.9%, n=103) reported they were likely to continue to share feedback through those channels.
@@ 4.2.3. Cultural and Institutional Legitimacy
Interactive feedback features may appeal more to users with certain cultural and institutional backgrounds, which can influence who submits feedback and the types of feedback they provide. Indeed, most of our survey respondents were young and identified as men, which reflects ChatGPT’s reported user base (Vogels, 2023) and other studies of contributors to human feedback datasets (Kirk et al., 2024). Generally, feedback types were proportional across demographic groups, though we found that respondents’ region had a moderate association with providing open-text feedback ($\chi^{2}$ = 13.990, adj. $p$-value = 0.023, df = 3, Cramér’s V = 0.192).
Instead, our survey results suggest that feedback features appeal more to users with frequent engagement or familiarity with AI, ML, and data science. Frequent ChatGPT users, particularly daily users, were more likely to provide thumbs-up/down feedback ($\chi^{2}$ = 13.689, adj. $p$-value = 0.045, df = 4, Cramér’s V = 0.166) than would be expected if there was no relationship between the variables, and those who previously used AI tools ($\chi^{2}$ = 8.926, adj. $p$-value = 0.023 df = 1, Cramér’s V = 0.154) were more likely to provide open-text feedback than would be expected. We also observed moderate associations suggesting more frequent ChatGPT users and those often engaging with AI-related topics were more likely to submit open-text feedback than would be expected if there was no relationship between these variables. However, these associations did not reach statistical significance after adjusting $p$-values. Additionally, we noted a borderline significant association between education or training in AI and open-text feedback submission (adj. $p$-value = 0.059).
4.3. Summary of Affordances
Table 1 outlines the affordances of the feedback features in the ChatGPT interface based on the analysis above, organised by the six mechanisms proposed by the framework. The affordances are linked to the conditions (perception, dexterity, and cultural and institutional legitimacy) that shape how users interact with and perceive the feedback features.
@@ 4.3. Summary of Affordances
Mechanism Affordance Condition
Request Feedback formatted in the ‘helpful, honest, and harmless’ framework Feedback from users to developers (Dexterity)
Demand Interaction with ChatGPT before feedback submission Status as a user (Dexterity)
Encourage Thumbs-up/down feedback; Feedback after regenerating a response; Performance-related open-text feedback Ease of use and accessibility of interface (Perception); Frequency of use and familiarity with the system (Perception, Dexterity); AI/ML experience and knowledge (Cultural and institutional legitimacy)
Discourage Feedback from collectives; Discussions among users Individualised feedback design (Cultural and institutional legitimacy)
Refuse Modifying feedback over time Stochastic nature of the system (Dexterity)
Allow Alternative channels for feedback and discussion (e.g., social media, online forums) Availability of alternative channels (Dexterity); Accessibility and appeal of feedback features to certain user groups (Cultural and institutional legitimacy)
@@ 5. Discussion
Participation is never fully open or anarchic; it is always formatted with “specific rules and scripts” (Kelty, 2020). Feedback features in interfaces to LLMs format user input into RLHF workflows. Therefore, the design of these features shapes the legitimacy of user participation in the iteration of LLMs. Our analysis of feedback features in ChatGPT suggests that the feedback format imposed by the interface is simple, individualised, unidirectional, and focused on performance-related concerns. This format significantly constrains user participation and risks becoming self-perpetuating—users may internalise limited roles in the iteration of LLMs, discouraging them from seeking broader engagement or viewing themselves as co-creators in LLM development. This phenomenon is not unique to LLMs; similar patterns have been observed in other technological domains, such as social media platforms (Gillespie, 2020).
Affordance theory as a design tool requires an orientation that “technologies, and the worlds they build, need not be as they are and can, instead, be otherwise” (Davis, 2023). As LLMs increasingly become infrastructure, it is important to consider participation beyond mere preference elicitation. In this section, we discuss two directions for the (re)design of feedback processes. Initially, we review recent developments in deliberative approaches to feedback, such as recent democratic experiments of OpenAI and Anthropic. However, we argue that users and stakeholders can never get to the heart of closed systems by commenting on model outputs. Instead, we encourage a shift in approach from alignment of model outputs to specific user preferences to dialogue with the public about the purpose and applications of LLMs, and attention to the ongoing infrastructuring work required to support and sustain diverse ‘publics’ over time.
@@ 5.1. Deliberative Processes
As previously noted, the existing feedback features request feedback from users to developers and discourage both feedback from groups of users and discussion amongst users. In other words, feedback through existing features is atomised and unidirectional (Wang et al., 2023). Additionally, as the number of ChatGPT users has grown and the system has stabilised over time, user input has been constrained from a wide to a narrow aperture—from free-text feedback forms to open-text feedback windows to singular buttons focused on crowdsourcing inputs to RLHF. These latest feedback formats are inadequate for resolving social and ethical considerations, requiring us to consider other design approaches.
@@
If algorithmic fairness is primarily a political task, as Wong (Wong, 2020) argues, rather than solely a technical task, we must consider how to resolve issues of bias and fairness politically rather than technically. Alongside aggregating citizen preferences through voting, deliberation among citizens is a key element of democratic legitimacy (Fishkin, 1991; Knight and Johnson, 1994; Curato et al., 2017). Deliberative democracy involves an association of members who govern their affairs by deliberation among the members (Elster and Przeworski, 1998)—making collective decisions determined by those subject to the decisions: not only their preferences but also their reasoning (Cohen, 2005). Such associations tend to be between broad groups of people—in this context, users, developers, and potentially groups of both—and are directed towards participants considering others’ perspectives and potentially changing their own over time.
@@
Interestingly, survey respondents identified several other platforms on which they discuss their experiences and preferences regarding ChatGPT, engaging in the kind of collective deliberation that the existing feedback features discourage. For example, OpenAI discussion forums allow communication between users and developers and subreddits such as r/ChatGPT allow discussion amongst users (without monitoring or oversight by OpenAI). However, it is unclear whether or how these discussions are input into development processes; while they function as evaluative feedback, there is no clear cybernetic feedback loop.
@@
Some technology companies have been experimenting with deliberative modes for feedback about AI systems. Recent efforts by OpenAI to fund research into technical solutions for “democratic inputs into AI” (Eloundou and Lee, 2024) and experiments by Anthropic to draft a constitution for Claude through deliberative processes (Huang et al., 2024) have attracted significant attention. These efforts go beyond eliciting and aggregating user preferences towards encouraging deliberation amongst people with diverse perspectives (including non-users) and demonstrating how the preferences and reasoning of multiple participants are input (as cybernetic feedback) into development workflows. In addition, Meta has experimented with Community Forums to deliberate on the rulesets for Generative AI (Broxmeyer, 2024), and Google has developed the STELA process for deliberative discussions with historically underrepresented groups in the U.S.A. to understand their priorities and concerns with respect to model outputs (Bergman et al., 2024).
@@
While these processes address some of the limitations of existing feedback features, they remain largely experimental and peripheral to the dominant method of feedback submission through browser-based interfaces. Most of these approaches (with the Google example as an exception) take as a given that consensus on the output of the process is desirable and achievable. However, refusal, agonism and dissensus are key elements of democratic processes (DiSalvo, 2015, 2010; Zong and Matias, 2024). When consulting multiple and diverse communities, values and preferences for AI design may be in conflict (Sambasivan et al., 2021; Qadri et al., 2023) and difficult to reconcile (Santurkar et al., 2023; Durmus et al., 2023; Miller et al., 2007). This is particularly the case where the processes are managed by those with commercial interests in consensus outcomes, reflecting long-standing concerns in participatory design (Sanders and Stappers, 2014).
@@
Moreover, such democratic inputs to the alignment problem, much like the existing feedback processes they are intended to augment or replace, narrowly scope participation around the output of specific families of models rather than engaging participants in earlier stages of the development process. Although deliberative processes may address some limitations of existing mechanisms and change how companies like OpenAI and Anthropic make decisions about model behaviour, they do not significantly alter the power dynamics between users, the public, and platforms. Users and other participants remain as consumers rather than co-creators, limited in their influence over core aspects of earlier stages of development.
@@ 5.2. Infrastructuring Participation
Our study reveals a tension between technology companies’ stated intentions and their implementation of feedback features in LLMs. ChatGPT exemplifies this disparity: what OpenAI initially sought (diverse, substantive feedback to uncover potential risks and harms) gave way to what they could effectively use within their existing workflows (easily processed, formatted data). This change not only constrains user input but also, as our survey indicates, favours frequent ChatGPT users and those with AI experience, narrowing the group of users influencing LLM iteration.
While eliciting user preferences as input into RLHF workflows may be well-intentioned and useful for technology companies to improve model performance, it is inadequate for the larger task facing us all. We all face the challenge of publicly, deliberatively, and inclusively addressing the ongoing tectonic shifts in information and knowledge generation across economic, social, cultural, and political domains before LLMs become unremarkable and unquestionable infrastructure.
If we are to shift our focus from aligning model outputs to specific user preferences towards dialogue with the public about the purpose and applications of LLMs, then we must think critically about who will form this ‘public’, how it will form, and what channels facilitate its participation. Currently, technology companies configure participation from atomised users of dialogue agents like ChatGPT, tightly defining processes for feedback on model outputs. In ChatGPT’s case, the ‘public’ is constructed primarily to address a matter of concern to OpenAI rather than those engaged as participants. Moreover, groups or institutions affected by the development and deployment of LLMs lack direct mechanisms to influence these feedback processes (Wang et al., 2023).
Technology companies might argue for increasing the number and diversity of participants providing feedback through existing channels. However, this approach misses the core issue. Drawing parallels with Arnstein’s (Arnstein, 1969) critique of participation in the context of urban planning, we argue that identifying problems in the proceduralisation of participation through feedback features in LLMs is not a call for more involvement but for a fundamental rethinking of the process. We echo Wang et al. (Wang et al., 2023) in urging companies to focus on the participation of groups or institutions representing populations likely to be significantly impacted by a given LLM. Infrastructuring such participation requires deep engagement with the methodological tools of Participatory Design (PD), which has a long-standing practice of critical thinking about how to form and sustain ‘publics’ to address participants’ concerns over time (DiSalvo, 2009; Le Dantec and DiSalvo, 2013). When implemented effectively, PD prioritises future issues concerning participants rather than solely focusing on the immediate concerns of designers or researchers.
@@
Recent research offers promising directions for rethinking processes and infrastructure for participation in LLM development. The STELA process (Bergman et al., 2024) demonstrates one method for conducting community-centred deliberative discussions about LLMs, albeit focused on aligning models with the preferences of specific groups of people. Suresh et al. (Suresh et al., 2024) expand on this concept by proposing a comprehensive framework for domain-specific community infrastructure, including shared technical resources, norms, and governance structures. Their approach aims to enable more meaningful participation by groups and institutions in shaping the development and application of foundation models in specific domains. Building on these ideas, we encourage participatory workshops with stakeholders in fields like education, healthcare, and journalism as a practical mechanism for realising and refining such domain-specific infrastructure. However, implementing these participatory approaches faces significant challenges: maintaining local relevance while extending across domains and regions, translating insights into actionable development inputs, and securing sustainable funding for ongoing infrastructure in the rapidly changing landscape of AI development. To advance meaningful public participation in the iteration of LLMs, the focus of research and effort must be on addressing these challenges rather than defaulting to procedural and perfunctory engagement.

### v2

@@ 4.2. Conditions
The mechanisms described above operate through three conditions: perception, dexterity, and cultural and institutional legitimacy. Each condition modifies how mechanisms function for different users, creating variation in how the interface affords participation.
@@ 4.2.1. Perception
Respondents generally perceived the feedback features in the interface as accessible and user-friendly. As shown in Table 5, over three-quarters reported confidence in their ability to evaluate and submit feedback about the performance and/or impact of generative AI systems such as ChatGPT. Of those who submitted at least one of the three feedback modes, almost all found doing so through the features extremely or somewhat easy (95.1% nn=442).
Table 5. Confidence Submitting Feedback (nn=524)
Confidence level Percentage
Strongly agree 21.8%
Somewhat agree 56.5%
Neither agree nor disagree 15.1%
Somewhat disagree 5.7%
Strongly disagree 1.0%
Total 100.0%
However, some users were more confident than others in submitting feedback. Those who frequently engaged with AI-related topics expressed more confidence than would be expected if there was no relationship between these factors and confidence (weak association; see Table 7 for all statistically significant relationships). Conversely, users with less frequent engagement were less confident than would be expected. Younger users (18-34) also showed a weak association with higher confidence in the initial analysis, though this finding was sensitive to the statistical approach used; the association lost significance when tested against the original five-point scale.
The differences in perception impact which mechanisms operate for different user groups. For confident users familiar with AI topics, the interface appears to encourage feedback submission through its accessible design. For the 21.8% of respondents who did not express confidence in evaluating and submitting feedback (those responding “Neither Agree nor Disagree” to “Strongly Disagree”), these same interface features may effectively discourage participation despite their availability.
@@ 4.2.2. Dexterity
Most respondents reported no barriers to accessing ChatGPT or submitting feedback (77.3%, nn=519), suggesting that users with varying levels of dexterity can provide input. Among those who encountered barriers, the most common were language and internet connectivity issues. We did not find significant associations between most demographic factors or AI experience and knowledge factors and those who experienced barriers. While those working in AI were more likely to report experiencing barriers than would be expected if there was no relationship between the variables, the association was weak (see Table 7).
However, as noted earlier, the requirement to prompt the dialogue agent before submitting feedback effectively demands system usage as a prerequisite for participation and refuses feedback from non-users. Our sample, therefore, does not include non-user perspectives, and given that most respondents used ChatGPT at least weekly, it also includes limited input from infrequent users. Notably, of those who completed the survey but did not indicate they had submitted feedback through one of the three options, the most common reason was that they were unaware of the feedback options (see Table 6). For these users, what designers intended to allow is functionally refused.
Table 6. Main Reasons for Not Submitting Feedback (nn=78)
Reason Count Percentage
Unaware of feedback option 27 34.6%
Not enough time or interest in submitting feedback 21 26.9%
Satisfied with performance and have not felt need to submit feedback 17 21.8%
Unsure how to provide feedback 7 9.0%
Other 5 6.4%
Unsatisfied with performance, but do not believe feedback will lead to improvements 1 1.3%
Total 78 100.0%
Also, as noted earlier, some respondents found alternative channels and platforms (outside the ChatGPT interface) to express their opinions and participate in discussions about the system. These results suggest that while the interface discourages certain forms of participation (e.g., collective feedback or discussions among users), users with sufficient dexterity sought platforms that allow forms of engagement the original interface refuses.
@@ 4.2.3. Cultural and Institutional Legitimacy
Feedback features may appeal more to users with certain cultural and institutional backgrounds, which can influence who submits feedback and the types of feedback they provide. Indeed, most of our survey respondents were young and identified as men, consistent with ChatGPT’s reported user base (Vogels, 2023) and studies of contributors to human feedback datasets (Kirk et al., 2024). Generally, feedback types were proportional across demographic groups, though geographic region moderately correlated with open-text feedback submission (see Table 7).
More significantly, the survey insights in Table 7 suggest that feedback features appeal more to users with frequent engagement or familiarity with AI, ML, and data science. Frequent ChatGPT users, particularly daily users, were more likely to provide thumbs-up/down feedback and those who previously used AI tools were more likely to provide open-text feedback than would be expected if there was no relationship between the variables. We also observed that education in AI correlated with open-text feedback submission, though this relationship fell just short of statistical significance (adj. $p$-value = 0.059, see Table 38). These patterns suggest that technical background, rather than demographic characteristics, most strongly influenced how our respondents engaged with feedback features.
These insights suggest how cultural and institutional legitimacy conditions shape the operation of affordance mechanisms. While the interface allows feedback from any user, it appears to more successfully encourage feedback from those whose technical knowledge or experience aligns with existing institutional norms in AI development.
Table 7. Significant Associations Between User Characteristics and Feedback Behaviours
Relationship $\chi^{2}$ $p$-value Cramér’s V Interpretation
Thumbs-Up/down and After Regenerating 50.419 ¡.001 .319 Moderate
Education/Training in AI and Confidence 12.773 .023 .160 Weak
Working in AI and Barriers 9.821 .023 .143 Weak
Region and Open-text feedback 13.990 .023 .192 Moderate
Prior AI use and Open-text feedback 8.926 .023 .154 Weak
Thumbs-Up/down and ChatGPT frequency 13.689 .045 .166 Weak
AI topics frequency and Confidence 20.726 .045 .144 Weak
Age and Confidence 17.068 .045 .131 Weak
@@
Feature(s) Mechanism of Affordance Outcomes For Whom/Condition
Checkboxes and structure in feedback windows Request Feedback formatted in the “helpful, honest, and harmless” framework Feedback from users to developers (Dexterity)
Login and interaction requirement Demand System usage as prerequisite for participation Excludes non-users (Dexterity)
Thumbs-up/down buttons; Regeneration feedback options; Open-text box with categorical selection options Encourage Simple, frequent feedback; Performance-focused feedback; Technical issue reporting For users with AI/ML experience and knowledge (Cultural and institutional legitimacy); For frequent users familiar with the system (Perception, Dexterity)
Single-user login Discourage Collective feedback; Discussion among users Excludes collectives of users or stakeholders (Cultural and institutional legitimacy)
No edit functionality after submission; Stochasticity of LLMs Refuse Modification of feedback over time; Comparative evaluation of responses For all users (Dexterity)
Access to external platforms Allow Alternative channels for feedback and discussion For technically knowledgeable users (Dexterity); For users with social media presence (Cultural and institutional legitimacy)
4.3. Synthesis: How Affordances Shape Participation
Table 8 outlines the affordances of the feedback features in the ChatGPT interface based on the analysis above, organised by the six mechanisms proposed by the framework. The affordances are linked to the conditions that shape how users interact with and perceive the feedback features. In summary, the interface requests feedback formatted in the “helpful, honest, and harmless” framework, demands interaction before feedback submission, and encourages simple, frequent feedback focused on performance improvement. However, the interface discourages collective feedback and discussions among users, while allowing alternative channels for such discussions, and refuses feedback modification over time. Additionally, our survey indicates that the feedback features may appeal more to users with frequent engagement, prior experience, and knowledge of AI, ML or data science.
@@ 5. Discussion
Participation is never fully open or anarchic; it is always formatted with “specific rules and scripts” (Kelty, 2020). Feedback features in interfaces to LLMs format user input into RLHF workflows, which both legitimises user participation in the iteration of LLMs and constrains such participation to unidirectional channelling of atomised, de-contextualised user preferences. The interface design also channels user input toward easily processed feedback focused on performance-related concerns; 76% of our respondents used thumbs-up/down feedback, but only 29% provided open-text feedback, with technical or performance motivations significantly outweighing social or ethical concerns. This highlights a clear tension between technology companies’ stated intentions (e.g., OpenAI initially sought diverse, substantive feedback to uncover potential risks and harms (OpenAI, 2022)) and how these features are actually used (simple, performance-focused feedback). While ostensibly democratising participation, we ask whether participation through feedback features is in name only, part of a “proceduralized participation… [that] can often be implemented without any challenge to existing power relations?” (Kelty, 2020).
Our sample primarily comprises young, tech-savvy early adopters of ChatGPT. While this limits generalisability, the demographic skew itself highlights how affordances of feedback features in initial deployments of LLMs might privilege certain groups of users, and the narrow forms of participation we observed among these users demonstrate how the interface’s affordances encouraged simple, performance-focused feedback in this context.
These constraints among early users raise broader questions about how such feedback processes might function as systems scale. If early adopters primarily provide simple, performance-focused feedback due to the interface design, users with less technical experience and knowledge, or facing other barriers, might find it even more difficult to articulate complex social or ethical concerns through these channels. This indicates the potential for a self-perpetuating dynamic where users internalise limited roles as participants in LLM iteration. This phenomenon is not unique to LLMs; similar patterns have been observed in other technological domains, such as social mediaplatforms (Gillespie, 2020).
Affordance theory as a design tool requires an orientation that “technologies, and the worlds they build, need not be as they are and can, instead, be otherwise” (Davis, 2023). Drawing on affordance theory, we discuss two potential directions for the (re)design of feedback processes to move beyond the constraints revealed by our analysis, aiming at shifting participation from procedural to more substantive forms.
@@ 5.1. Deliberative Processes
As discussed in Section 2.1, resolving social and ethical issues in AI deployment is increasingly framed as a political task rather than a technical task. One potential direction for redesigning feedback processes to address such issues politically is to incorporate elements of deliberation—structured dialogue among users and (potentially) with developers.
@@
Alongside aggregating citizen preferences through voting, deliberation among citizens is a key element of democratic legitimacy (Fishkin, 1991; Knight and Johnson, 1994; Curato et al., 2017). Deliberative democracy involves an association of members who govern their affairs through deliberation among the members (Elster and Przeworski, 1998)—making collective decisions based not only on their preferences but also on their reasoning (Cohen, 2005). Such associations typically involve broad groups of people and encourage participants to consider others’ perspectives and potentially revise their own views over time
@@
Deliberative processes could change the existing affordances of user participation in iterating LLMs. For example, where current interfaces discourage reasoned discussion among users, deliberative platforms could explicitly request reasoning from participants and allow for bidirectional communication between users and developers, along with evolving perspectives, moving toward genuine feedback loops rather than the unidirectional channels currently prevalent. Interestingly, some respondents identified several other platforms on which they discuss their experiences and preferences regarding ChatGPT, engaging in the kind of collective deliberation that the existing feedback features discourage. For example, OpenAI discussion forums allow communication between users and developers and subreddits such as r/ChatGPT allow discussion among users (without monitoring or oversight by OpenAI). However, it is unclear whether or how these discussions feed back into development.
@@
Some technology companies have been experimenting with deliberative modes for feedback about AI systems. Recent efforts by OpenAI to fund research into technical solutions for “democratic inputs into AI” (Eloundou and Lee, 2024) and experiments by Anthropic to draft a constitution for Claude through deliberative processes (Huang et al., 2024) have attracted significant attention. These efforts go beyond eliciting and aggregating user preferences. They encourage deliberation among users and (potentially) other stakeholders, while demonstrating how multiple participants’ preferences and reasoning feed into development workflows. In addition, Meta has experimented with Community Forums to deliberate on the rulesets for Generative AI (Broxmeyer, 2024), and Google has developed the STELA process for deliberative discussions with historically underrepresented groups in the U.S.A. to understand their priorities and concerns with respect to model outputs (Bergman et al., 2024).
@@
While these processes address some of the limitations of existing feedback features, they remain largely experimental and peripheral to the dominant method of feedback submission through browser-based interfaces. Most of these approaches (with the STELA process as an exception) take as a given that consensus on the output of the process is desirable and achievable. However, refusal, agonism and dissensus are all key elements of democratic processes (DiSalvo, 2015, 2010; Zong and Matias, 2024). When consulting multiple and diverse communities, values and preferences for AI design may be in conflict (Sambasivan et al., 2021; Qadri et al., 2023) and difficult to reconcile (Santurkar et al., 2023; Durmus et al., 2024; Miller et al., 2007). Such difficulty is particularly acute where the processes are managed by those with commercial interests in consensus outcomes, reflecting long-standing concerns in participatory design (Sanders and Stappers, 2014).
@@
Moreover, such democratic inputs to the alignment problem, much like the existing feedback features they are intended to augment or replace, narrowly scope participation around the output of specific families of models, refusing participants meaningful engagement with earlier development stages (Delgado et al., 2023). Although potentially addressing some limitations of current feedback features, users and stakeholders remain as consumers rather than co-creators.
@@ 5.2. Infrastructuring Participation
While our insights are based on early adopters, the constraints revealed by our analysis highlight the limitations of current processes for feedback on harmful model outputs or risks from LLM deployment. Existing processes appear to be efficient channels for atomised preferences from specific user groups, but insufficient conduits for dialogue on social and ethical issues regarding LLMs. Echoing Arnstein’s (Arnstein, 1969) critique of participation in the context of urban planning, identifying problems in the proceduralisation of participation through existing processes is not simply a call for more involvement, but instead requires a fundamental rethinking of the process.
We advocate for moves beyond aligning model outputs to specific user preferences toward feedback processes that embrace productive disagreement and plurality. As DiSalvo (DiSalvo, 2010) argues, designers can enhance democratic engagement by “creating and enabling … spaces of contest” where “difference and dissensus are brought forward and the assumptions and actions that shape power relations and influence are revealed and challenged” (DiSalvo, 2010). To create feedback channels capable of supporting these more contested forms of engagement requires actively “infrastructuring” participation—creating and sustaining social, technical, and institutional structures with different affordances to address matters of concern to participants (Le Dantec and DiSalvo, 2013). Such infrastructure must aim at changing the conditions of participation, to actively encourage engagement from stakeholders affected by LLM deployments, including non-users and those without technical backgrounds, and allow forms of input that address participants’ concerns beyond immediate systemperformance.
@@
Recent research offers promising practical directions. The STELA process (Bergman et al., 2024) demonstrates one method for conducting community-centred deliberative discussions about LLMs with marginalised groups who may or may not be users of LLMs, albeit focused on aligning models with the preferences of those groups of people. Suresh et al. (2024) expand on this concept by proposing a comprehensive framework for domain-specific community infrastructure, including shared technical resources, norms, and governance structures. Their approach aims at enabling participation by stakeholder groups and institutions to shape the development and application of foundation models in specific domains. Building on these ideas, we encourage participatory design workshops with stakeholders, in fields like education, healthcare, and journalism, that use techniques encouraging agonism (see, e.g., (Björgvinsson et al., 2012)) to surface conflicts and refine such domain-specific infrastructure.
Building and sustaining this infrastructure is not a one-off task; it requires ongoing effort to secure resources, build trust with participants, navigate conflict, ensure local relevance while extending across domains and regions, and translate outputs into actionable development inputs (Le Dantec and DiSalvo, 2013). However, these are not reasons to default to simpler, constrained modes of participation. If the goal of feedback processes is to afford meaningful public influence over technologies with profound societal impact, then the focus of research and effort must be on addressing such challenges to infrastructuring sustained, pluralistic, and sometimes contentious dialogue with affected stakeholders.